# thanks: ORCID ID: 0009-0003-9679-035Xthanks: ORCID ID: $0$-0001-7989-9271

### v1

@@ 4.3 Long-time drift and random motion
In the very small regions between the stable sections in Fig. 10(b) we observe shuttling with unequal numbers of loops around the left- and right-traveling states. This induces a long-time effective drift of the droplet [Fig. 11(a), Supplemental Video 7]. Corresponding asymmetric periodic orbits also emerge in the context of Shilnikov bifurcations with symmetry [80]. Further, for driving forces that correspond to a large number of loops ($l>30$) we also find shuttling motion, where the number of loops changes irregularly with time. This results in what appears to be long-time random drop motion [Fig. 11(b), Supplemental Video 8]. Because the relevant parameter regions for both cases are very small, here we do not discuss them any further.
Lastly, we remark that highly similar bifurcation cascades, where the Shilnikov condition holds, have been reported for systems of ordinary differential equations [84, 85, 78]. In Refs. [84, 85] they were also named “incomplete homoclinic scenarios” due to the apparent absence of a Shilnikov homoclinic orbit. Nevertheless, for the Koper model [78] in particular, the existence of such a homoclinic orbit was recently demonstrated [86].
@@ 4.4 Organization around higher-codimension bifurcations
Next, we examine how the foregoing bifurcations change when a second parameter, the diffusion constant $D\!=\!D_{1}\!=\!D_{2}$, is varied. The loci of the bifurcations discussed above are tracked in the $\left(\mu_{2},D\right)$-plane and presented in Fig. 5(c). The locus of the primary bifurcation DP1 from the base branch shows that the uniformly covered resting droplet is linearly stable for most of the shown parameter region. The Hopf bifurcations H+ and H- interchange positions at C±, rendering the simple traveling droplets unstable for all values of $\mu_{2}$. Of particular interest are the loci of the gluing bifurcation G and of the inferred Shilnikov bifurcation S. [^7] For a given $D$, we determine $\mu^{G}$ from the divergence of the period at the gluing bifurcation. We approximate $\mu^{S}$ by the location of the $l\!=\!8$-branch. When $D$ is increased, eventually the distance of the loci of G and S becomes very small [Fig. 5(c), region T]. Beyond this region, the one-parameter bifurcation diagram at fixed $D$ transforms drastically and only two branches of shuttling states exist with either $l\!=\!0$ or $l\!=\!2$ loops, forming a continuous parameter region in which shuttling is observed [Fig. 5(b)]. Notably, the period remains finite. It is instructive to consider the representation of states in the ($\theta_{l},\theta_{r}$)-plane in the region T [Fig. 8(c)]. Typical trajectories, e.g., near the gluing bifurcation additionally show several loops, indicating that the homoclinic connections at the gluing bifurcation also closely pass by the traveling states [Fig. 9(e)]. We can then expect that there exists a point in the ($\mu_{2},D$)-plane where the homoclinic connections collide with the traveling states, or equivalently, where the heteroclinic connections at the Shilnikov bifurcation meet at the central one-peak resting state. Then, the central resting state is simultaneously connected to both traveling states in a heteroclinic loop [Fig. 9(f)]. The corresponding codimension-2 point where G and S terminate is called a T-point [87, 88]. The two-parameter bifurcation diagram near this point is expected to be similar to Fig. 5(i), where the loci of G spiral into the T-point, while S approaches it in a straight line. [^8] We note that S is located in close proximity of the line of Hopf bifurcations H-, beyond which the simple traveling states are linearly stable. It is possible that the true S crosses the line of H- at a codimension-2 Shilnikov-Hopf point, and that when a third parameter is varied, the T-point crosses H- at a codimension-3 T-point-Hopf bifurcation. In these cases, homo- and heteroclinic connections to the traveling states transform into connections to the time-periodic states that emerge subcritically at H-. The one-parameter snaking structure of the shuttling states and the two-parameter spiral of G remain qualitatively similar when crossing these bifurcations [89, 90]. We do not resolve this structure but note that in region T we find linearly stable traveling, shuttling, crawling and uniformly covered resting droplets. Lastly, we observe that the shuttling states that emerge from S cannot be observed for low $D$, corresponding to an end of the bifurcation line [region A in Fig. 5(c)]. Because the Shilnikov-condition $|\rho/\lambda_{3}|<1$ is still fulfilled in this region, we speculate that here, the line S turns around toward greater $D$. For a discussion of other possible termination mechanisms, see Ref. [91].
@@
Up to here, we have discussed 2D droplets on 1D substrates. However, our considerations of the feedback loop in Sec. 3 also apply to the case of a 3D drop on a 2D substrate. The liquid transport flux $\bm{j}_{h}$ and the surfactant transport fluxes $\bm{j}_{1}$ and $\bm{j}_{2}$ in Eqs. (7) are then two-dimensional and can exhibit, e.g., vortex structures. This may result in more complex dynamics, particularly in parameter regions where 2D drops show shuttling or crawling motion. The snapshots from a time simulation presented in Fig. 12 show a typical change of the direction of droplet motion (Supplemental Video 9). Localized surface tension gradients are simultaneously excited in different regions of the droplet, breaking rotational symmetry and leading to complex liquid flow with several vortices near areas of increased surface tension. The Marangoni stresses between such regions then cause fluid flows that result in an effective attraction between neighbouring surface tension peaks. Ultimately, this results in an irregular dynamics of generation, attraction and merging of these localized structures. Droplets then no longer move along straight lines but can explore the entire substrate in a random walk. This is mediated by the repeated nucleation of protruding regions of high surface tension at various positions along the contact line. A typical trajectory is shown in Fig. 13(a). We characterize this motion in Fig. 13(b) by computing the mean squared displacement $\langle(\Delta r)^{2}\rangle(\tau)=\frac{1}{t_{B}-t_{A}}\int_{t_{A}}^{t_{B}}||\bm{r}(t+\tau)-\bm{r}(t)||\>\text{d}t$, where $t_{A}$ and $t_{B}$ are the start and end times of the trajectory (Appendix E). We find two scaling regimes characterized by different power laws $\langle(\Delta r)^{2}\rangle\sim\tau^{\alpha}$. At short times, we have $\alpha\approx 2$, i.e., the motion is ballistic. At long times, we have $\alpha\approx 9/10$ and droplet motion is to a good approximation diffusive ($\alpha=1$). Note that the small undulations visible in this regime correspond to decaying oscillations in the velocity auto-correlation function $\langle\bm{v}\cdot\bm{v}\rangle(\tau)=\frac{1}{2}\frac{\text{d}^{2}\langle(\Delta r)^{2}\rangle}{\text{d}\tau^{2}}$ [92]. This implies that droplets are slightly more likely to reverse their direction of travel than to turn in any other direction (for pure shuttling motion one would have a non-decaying oscillation of $\langle\bm{v}\cdot\bm{v}\rangle$). Therefore, the random walk is not entirely uncorrelated. Interestingly, at small liquid volumes shuttling motion is recovered (not shown).
@@
Figure 12: Snapshots from a time simulation of a large 3D droplet performing a random walk. The columns show (a) the film thickness, (b) the surface tension, (c) the streamlines of the liquid transport flux $\bm{j}_{h}$ and (d) both the streamlines of the total surfactant transport flux $\bm{j}_{1}+\bm{j}_{2}$ (white lines) and the local total source term of surfactant $\mathcal{B}_{1}+\mathcal{B}_{2}$ (color map). In (b), regions of high surface tension are colored in red. The individual rows are at different times $t_{0},\ldots,t_{3}$ and show a typical change in the direction of motion resulting from the merging of the existing surface tension peaks ($t_{0}$ and $t_{1}$) and a subsequent generation of new peaks at different positions ($t_{2}$ and $t_{3}$). Note that the streamlines in (c) converge in regions of low $\mathcal{B}_{1}+\mathcal{B}_{2}$ that are surfactant sinks. Arrows indicate the direction of motion. The parameters are $W=10,r=0.4,D_{1}=D_{2}=0.2,\beta_{1}=2,\beta_{2}=0.01,\mu_{1}=-1.4,\mu_{2}=4.175,\delta=1$ with a mean film thickness of $\bar{h}=3$. The periodic computational domain is $\left[0,100\right]\times\left[0,100\right]$. See also Supplemental Video 9.
@@
Figure 13: (a) Trajectory of the center of mass of a droplet undergoing a random walk from point $A$ to point $B$ (red dots) over an approximate time interval of $t_{B}\!-\!t_{A}\!=3\!\times\!10^{4}$. (b) Corresponding mean square displacement $\langle(\Delta r)^{2}\rangle$ for the trajectory in $(a)$. There are two scaling regimes with $\langle(\Delta r)^{2}\rangle\sim\tau^{\alpha}$. For short times, droplet motion is ballistic with $\alpha\approx 2$ and for long times it is approximately diffusive with $\alpha\approx 9/10$. The parameters are as in Fig. 12.
@@ 5 Conclusion and Outlook
Many biomimetic and biological systems involve chemo-mechanical coupling with interacting chemical reactions and hydrodynamic transport resulting in complex spatio-temporal behavior. To study such interactions, we have employed a conceptually simple mesoscopic hydrodynamic model that captures the dynamics of sessile droplets of a partially wetting liquid covered by autocatalytically reacting surfactants. The droplets are supplied with chemical fuel by an external chemostat. Notably, our description is based on nonequilibrium thermodynamics, ensuring the existence of a thermodynamically consistent passive (“dead”) limit. Our study has focused on the self-propulsion of droplets that emerges for sufficiently large nonequilibrium driving. The underlying mechanism is a positive feedback loop between the solutal Marangoni effect and the local chemical reactions. Importantly, in contrast to the commonly treated drop-based microswimmers with simple conversion reactions [39, 40], the mechanism is not based on differences in physical properties of different surfactants. The considered hydrodynamic scenario is also distinct since we treat droplets on a substrate in a stress-free ambient medium instead of fully immersed drops in a bulk fluid that contributes to the dynamics. Relaxing the latter assumption, one could additionally consider flows in the surrounding medium, e.g., along the lines of [61].
@@
Besides the simple self-propelled drops that move uniformly at constant speed, we find that droplets on 1D substrates may also crawl (periodic stick-slip motion) or shuttle (periodic back-and-forth motion). These types of motility can be explained by the positive feedback between Marangoni fluxes and the local reactor. However, in the context of dynamical systems they can also be discussed as periodic orbits in an effective low-dimensional phase space, where the corresponding trajectories closely pass fixed points representing traveling and resting states. Using this representation, we have identified two scenarios involving global bifurcations for the transitions between crawling and shuttling states. Based on two-parameter representations of the loci of a gluing and a Shilnikov global bifurcation in the $(\mu_{2},D)$-plane, we have further speculated that both bifurcations terminate at a codimension-2 point (T-point). Various different linearly stable resting and motile droplets exist in the vicinity of this point. It therefore acts as an organizing center in parameter space. Similar motifs arise in the description of biological systems, e.g., in the study of electrical or mechanical signaling in resting cells [93, 94] where small parameter changes allow cells to switch between different modes of operation. For the actin cortex of motile cells, it has been suggested on the basis of mass-conserving reaction-diffusion models that the bistability of various dynamical states of the cortex can be traced to codimension-2 points [95, 96]. Here, we find the organization of highly complex states of motility around a codimension-2 point for a full, spatially extended dynamical model that couples various transport and reaction processes in the presence of interfaces, which to our knowledge has not yet been reported.
@@
In our study of shuttling and crawling states, we have only employed direct numerical simulations since numerical continuation of such spatially localized time-periodic slow-fast states is numerically exceptionally challenging and computationally expensive. In consequence, we have only been able to obtain partial bifurcation diagrams without the unstable branch sections of crawling and shuttling droplets. In the future, our results could be extended by studying the entire bifurcation structure, e.g., by employing spatio-temporally adaptive continuation methods. One could thereby gain information about local bifurcations of such states. For example, we expect that branches of crawling states terminate either at Hopf bifurcations of traveling drops or at drift-pitchfork bifurcations of resting oscillating drops. Further, it is known that direction-reversing (shuttling) states can emerge in reflection-symmetry breaking Hopf bifurcations [77, 97, 98]. To verify our hypothesis that the two-parameter bifurcation diagram is organized about a T-point, it is necessary to employ numerical continuation of the associated heteroclinic and homoclinic connections involving the resting and traveling states - a task beyond present numerical tools for partial differential equations known to us. Lastly, we have turned to 3D droplets that, when large enough, explore the entire available substrate in an approximately diffusive random walk due to irregular excitation, attraction and merging of localized surface tension gradients near the contact line region. The transition between such random walks and other motility modes shall be investigated in the future.
@@
We expect that the discussed positive feedback loop is relevant to biomimetic droplet systems that feature higher-order chemical reactions of surface active species [99, 44, 100, 101]. Such reactions may occur, e.g., in auto- or cross-catalytic mechanisms and the formation of micellular structures. Moreover, while the presented simple model does not reflect the complex biochemistry of real biological cells, it may capture a motility mechanism relevant under prebiotic conditions, i.e., for protocells, where autocatalytic reactions may represent a simple form of molecular self-replication. Furthermore, because common recent descriptions of biomolecular condensates within cells are based on a similar thermodynamic structure [23, 22, 62, 63], we believe that similar states and higher-codimension organizing centers may be of importance if one includes cross- or autocatalytic effects and considers nonlinear mobility functions. In our case, the latter result from advection but they may also arise, e.g., for diffusion in crowded environments [102]. Finally, we remark that one may lift our present restriction to the case where the underlying reaction system does not form any patterns. Then one needs to amend the mobilities and reaction rates, e.g., by adding film thickness-dependent cutoffs of the rates to maintain a passive adsorption layer representing a macroscopically dry and therefore passive substrate.

### v3

@@ 4.3 Long-time drift and random motion
In the very small regions between the stable sections in Fig. 10(b) we observe shuttling with unequal numbers of loops around the left- and right-traveling states. This induces a long-time effective drift of the droplet [Fig. 11(a), Supplemental Video 7]. Corresponding asymmetric periodic orbits also emerge in the context of Shilnikov bifurcations with symmetry [91]. Further, for driving forces that correspond to a large number of loops ($l>30$) we also find shuttling motion, where the number of loops changes irregularly with time. This results in what appears to be long-time random drop motion [Fig. 11(b), Supplemental Video 8]. This irregular direction-reversal dynamics is likely connected to the presence of the pair of (Shilnikov) heteroclinic connections between the right- and left-traveling states, but in other systems with reflection and translation symmetry it may also result from symmetry-increasing bifurcations of already existing chaotic attractors [95]. Because the relevant parameter regions for both cases are extremely small, here we do not discuss them further.
Lastly, we remark that quite similar bifurcation cascades, where the Shilnikov condition holds, have been reported for systems of ordinary differential equations [96, 97, 89]. In Refs. [96, 97] they were also named “incomplete homoclinic scenarios” due to the apparent absence of a Shilnikov homoclinic orbit. Nevertheless, in particular for the Koper model [89], the existence of such a homoclinic orbit was recently demonstrated [98].
@@ 4.4 Organization around higher-codimension bifurcations
Next, we examine how the foregoing bifurcations change when a second parameter, the diffusion constant $D\!=\!D_{1}\!=\!D_{2}$, is varied. The loci of the bifurcations discussed above are tracked in the $\left(\mu_{2},D\right)$-plane and presented in Fig. 5(c). The locus of the primary bifurcation DP1 from the base branch shows that the uniformly covered resting droplet is linearly stable for most of the shown parameter region. The Hopf bifurcations H+ and H- interchange positions at C±, rendering the simple traveling droplets unstable for all values of $\mu_{2}$. Of particular interest are the loci of the gluing bifurcation G and of the inferred Shilnikov bifurcation S. [^7] For a given $D$, we determine $\mu^{G}$ from the divergence of the period at the gluing bifurcation. We approximate $\mu^{S}$ by the location of the l=8l\!=\!8-branch. When $D$ is increased, eventually the distance of the loci of G and S becomes very small [Fig. 5(c), region T]. Beyond this region, the one-parameter bifurcation diagram at fixed $D$ transforms drastically and only two branches of shuttling states exist with either l=0l\!=\!0 or l=2l\!=\!2 loops, forming a continuous parameter region in which shuttling is observed [Fig. 5(b)]. Notably, the period remains finite. It is instructive to consider the representation of states in the ($\theta_{l},\theta_{r}$)-plane in the region T [Fig. 8(c)]. Typical trajectories, e.g., near the gluing bifurcation additionally show several loops, indicating that the homoclinic connections at the gluing bifurcation also closely pass by the traveling states [Fig. 9(e)]. We can then expect that there exists a point in the ($\mu_{2},D$)-plane where the homoclinic connections collide with the traveling states, or equivalently, where the heteroclinic connections at the Shilnikov bifurcation meet at the central one-peak resting state. Then, the central resting state is simultaneously connected to both traveling states in a heteroclinic loop [Fig. 9(f)]. The corresponding codimension-2 point where G and S terminate is called a T-point [99, 100]. The two-parameter bifurcation diagram near this point is expected to be similar to Fig. 9(i), where the loci of G spiral into the T-point, while S approaches it in a straight line. [^8] We note that S is located in close proximity of the line of Hopf bifurcations H-, beyond which the simple traveling states are linearly stable. It is possible that the true S crosses the line of H- at a codimension-2 Shilnikov-Hopf point, and that when a third parameter is varied, the T-point crosses H- at a codimension-3 T-point-Hopf bifurcation. In these cases, homo- and heteroclinic connections to the traveling states transform into connections to the time-periodic states that emerge subcritically at H-. The one-parameter snaking structure of the shuttling states and the two-parameter spiral of G remain qualitatively similar when crossing these bifurcations [101, 102]. We do not resolve this structure but note that in region T we find linearly stable traveling, shuttling, crawling and uniformly covered resting droplets. Lastly, we observe that the shuttling states that emerge from S cannot be observed for low $D$, corresponding to an end of the bifurcation line [region A in Fig. 5(c)]. Because the Shilnikov condition $|\rho/\lambda_{3}|<1$ is still fulfilled in this region, we speculate that here, the line S turns around toward greater $D$. For a discussion of other possible termination mechanisms, see Ref. [103].
@@
Up to here, we have discussed 2D droplets on 1D substrates. However, our considerations of the feedback loop in Sec. 3 also apply to the case of a 3D drop on a 2D substrate. The liquid transport flux $\bm{j}_{h}$ and the surfactant transport fluxes $\bm{j}_{1}$ and $\bm{j}_{2}$ in Eqs. (7) are then two-dimensional vector fields and can exhibit, e.g., vortex structures. This may result in more complex dynamics, particularly in parameter regions where 2D drops show shuttling or crawling motion. The snapshots from a time simulation presented in Fig. 12 show a typical change of the direction of droplet motion (Supplemental Video 9). Localized surface tension gradients are simultaneously excited in different regions of the droplet, breaking rotational symmetry and leading to complex liquid flow with several vortices near areas of increased surface tension. The Marangoni stresses between such regions then cause fluid flows that result in an effective attraction between neighboring surface tension peaks. Ultimately, this results in an irregular dynamics of generation, attraction and merging of these localized structures. Sufficiently large droplets then no longer move along straight lines but can explore the entire substrate in a random walk. This is mediated by the repeated nucleation of protruding regions of high surface tension at various positions along the contact line. Interestingly, at small liquid volumes shuttling motion is recovered (not shown).
[…]
@@
Figure 12: Snapshots from a time simulation of a large 3D droplet performing a random walk. The columns show (a) the film thickness, (b) the surface tension, (c) the streamlines of the liquid transport flux $\bm{j}_{h}$ and (d) both the streamlines of the total surfactant transport flux $\bm{j}_{1}+\bm{j}_{2}$ (white lines) and the local total source term of surfactant $\mathcal{B}_{1}+\mathcal{B}_{2}$ (color map). The streamline thickness corresponds to the magnitude of the flux. Dashed contour lines in (b), (c) and (d) show the outline of the droplet (given by $h=1.1$). In (b), regions of high surface tension are colored in red. The individual rows are at different times $t_{0},\ldots,t_{3}$ and show a typical change in the direction of motion resulting from the merging of the existing surface tension peaks ($t_{0}$ and $t_{1}$) and a subsequent generation of new peaks at different positions ($t_{2}$ and $t_{3}$). Note that the streamlines in (c) converge in regions of low $\mathcal{B}_{1}+\mathcal{B}_{2}$ that are surfactant sinks. Arrows in (a) indicate the direction of motion. The parameters are $W=10,r=0.4,D_{1}=D_{2}=0.2,\beta_{1}=2,\beta_{2}=0.01,\mu_{1}=-1.4,\mu_{2}=4.175,\delta=1$ with a mean film thickness of $\bar{h}=3$. The periodic computational domain is $\left[0,100\right]\times\left[0,100\right]$. See also Supplemental Video 9.
@@
Figure 13: (a) Trajectory of the center of mass of a droplet undergoing a random walk from point AA to point $B$ (red dots) over an approximate time interval of $t_{B}\!-\!t_{A}\!=3\!\times\!10^{4}$. (b) Corresponding mean square displacement $\langle(\Delta r)^{2}\rangle$ for the trajectory in (a). There are two scaling regimes with $\langle(\Delta r)^{2}\rangle\sim\tau^{\alpha}$. For short times, droplet motion is ballistic with $\alpha\approx 2$ and for longer times it is approximately diffusive with $\alpha\approx 9/10$. The parameters are as in Fig. 12.
@@ 5 Conclusion and Outlook
Many biomimetic and biological systems involve chemomechanical coupling with interacting chemical reactions and hydrodynamic transport resulting in complex spatio-temporal behavior. To study such interactions, we have employed a conceptually simple mesoscopic hydrodynamic model that captures the dynamics of sessile droplets of a partially wetting liquid covered by autocatalytically reacting surfactants. The droplets are supplied with chemical fuel by an external chemostat. Notably, our description is based on nonequilibrium thermodynamics, ensuring the existence of a thermodynamically consistent passive (“dead”) limit. Our study has focused on the self-propulsion of droplets that emerges for sufficiently large nonequilibrium driving. The underlying mechanism is a positive feedback loop between the solutal Marangoni effect and the local chemical reactions. Importantly, in contrast to the commonly treated drop-based microswimmers with simple conversion reactions [41, 42], the mechanism is not based on differences in physical properties of different surfactants. The considered hydrodynamic scenario is also distinct since we treat droplets on a substrate in a stress-free ambient medium instead of fully immersed drops in a bulk fluid that contributes to the dynamics. Relaxing the latter assumption, one could additionally consider flows in the surrounding medium, e.g., along the lines of [71]. Similarly, the present model may be augmented to account for the surfactant dynamics in the bath (which may correspond to either the droplet bulk or the ambient medium), i.e., by treating soluble surfactants [66] or surfactants diffusing like a vapor phase [105].
@@
Besides the simple self-propelled drops that move uniformly at constant speed, we find that droplets on 1D substrates may also crawl (periodic stick-slip motion) or shuttle (periodic back-and-forth motion). These types of motility can be explained by the positive feedback between Marangoni fluxes and the local reactor. However, in the context of dynamical systems they can also be discussed as periodic orbits in an effective low-dimensional phase space, where the corresponding trajectories closely pass fixed points representing traveling and resting states. Using this representation, we have identified two scenarios involving global bifurcations for the transitions between crawling and shuttling states. Based on two-parameter representations of the loci of a gluing and a Shilnikov global bifurcation in the $(\mu_{2},D)$-plane, we have further speculated that both bifurcations terminate at a codimension-2 point (T-point). Various different linearly stable resting and motile droplets exist in the vicinity of this point. It therefore acts as an organizing center in parameter space. Similar motifs arise in the description of biological systems, e.g., in the study of electrical or mechanical signaling in resting cells [106, 107] and stick-slip cycles of motile cells [108] where small parameter changes allow cells to switch between different modes of operation. For the actin cortex of motile cells, it has been suggested on the basis of mass-conserving reaction-diffusion models that the bistability of various dynamical states of the cortex can be traced to codimension-2 points [109, 110]. Here, we find the organization of highly complex states of motility around a codimension-2 point for a full, spatially extended dynamical model that couples various transport and reaction processes in the presence of interfaces, which to our knowledge has not yet been reported.
@@
In our study of shuttling and crawling states, we have only employed direct numerical simulations since numerical continuation of such spatially localized time-periodic slow-fast states is numerically exceptionally challenging and computationally expensive. In consequence, we have only been able to obtain partial bifurcation diagrams without the unstable branch sections of crawling and shuttling droplets. In the future, our results could be extended by studying the entire bifurcation structure, e.g., by employing spatio-temporally adaptive continuation methods. One could thereby gain information about local bifurcations of such states. For example, we expect that branches of crawling states terminate either at Hopf bifurcations of traveling drops or at drift-pitchfork bifurcations of resting oscillating drops. Further, it is known that direction-reversing (shuttling) states can emerge in reflection symmetry-breaking Hopf bifurcations [88, 111, 112]. To verify our hypothesis that the two-parameter bifurcation diagram is organized about a T-point, it is necessary to employ numerical continuation of the associated heteroclinic and homoclinic connections involving the resting and traveling states - a task beyond present numerical tools for partial differential equations known to us. Lastly, we have turned to 3D droplets that, when large enough, explore the entire available substrate in an approximately diffusive random walk due to irregular excitation, attraction and merging of localized spots of surface tension gradients near the contact line region. The transition between such random walks and other motility modes shall be investigated in the future. Here, we expect that particularly the droplet volume represents an important control parameter, since larger droplets can accommodate a larger number of localized surface tension peaks, resulting in more complex behavior. This holds for both 2D and 3D drops, i.e., we find that large 2D droplets ($\sim 50\%$ larger volume than the one used in Sections 4.1-4.4) also exhibit highly irregular dynamics. Furthermore, the linear stability of the traveling and resting states may change with the drop volume (e.g., the symmetric one-peak resting state may become linearly stable for small volumes), thus altering the bifurcation picture. In this regard, we expect that the results presented in Sec. 4 are valid for drops that can accommodate one or two surface tension peaks.
@@
The discussed positive feedback loop is likely relevant to many biomimetic droplet systems that feature higher-order chemical reactions of surface active species [113, 46, 114, 63, 64]. Such reactions may occur, e.g., in auto- or cross-catalytic mechanisms and the formation of micellar structures. Moreover, while the presented simple model does not reflect the complex biochemistry of real biological cells, it may capture a motility mechanism relevant under prebiotic conditions, i.e., for protocells, where autocatalytic reactions may represent a simple form of molecular self-replication. Additionally, the here studied forms of motility (crawling, shuttling with and without net drift, irregular motion) appear in models and experiments for amoeba [115] and Physarum droplets [116, 117, 14]. Although related theoretical descriptions may include mechanical effects that go beyond the present hydrodynamic model (e.g., viscoelasticity), it is worth pointing out that they typically feature a (linear) coupling of a nonlinear chemical kinetics and active mechanical stresses, in analogy to the here considered coupling of surface tension effects and autocatalytic chemical reactions via surfactants. Furthermore, because common recent descriptions of biomolecular condensates within cells are based on a similar thermodynamic structure [25, 24, 72, 73], we believe that also there similar states and higher-codimension organizing centers may be of importance if one includes cross- or autocatalytic effects and considers nonlinear mobility functions. In our case, the latter result from advection but they may also arise, e.g., for diffusion in crowded environments [118]. Finally, we remark that one may lift our present restriction to the case where the underlying reaction system does not form any patterns. Then one needs to amend the mobilities and reaction rates, e.g., by adding film thickness-dependent cutoffs of the rates to maintain a passive adsorption layer representing a macroscopically dry and therefore passive substrate.